• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name kelleher

    Log entry time 08:10:38 on April 27, 2007

    Entry number 200269

    keyword=Owl Shift Summary

    Friday, April 27 Owl Shift

    Plan: take beam

    00:00 Shift began during run 2878
    00:42 Problems with right S2m in the end of run files
    01:20 Rebooted the trigger supervisor crate (and CODA and ROCs)
    02:13 Download trigger
    03:09 raster tripped, MCC tried to reboot IOC, called Bill Gunning
    04:05 Bill Gunning controlled access to fix raster
    05:02 raster check (looks ok)
    05:06 resume production running
    06:46 target to dummy for background run
    07:34 move target back to 4He

    Run Summary

    Run Begin End Notes
    -----------------------------------
    2878 00:02 00:19
    2879 00:19 00:36
    2880 00:40 00:48
    2881-2883 Junk troubleshooting runs
    2884 01:18 01:36 Back to production, detector replay okay
    2885 01:37 01:52
    2886 01:53 02:12
    2887 02:15 02:38
    2888 02:38 02:57
    2889 02:57 03:09 raster tripped
    2890 junk
    2891 05:02 05:06 20uA raster scan
    2892 05:06 05:23 stopped after 10 min. w/o beam
    2893 05:25 05:39 beam restored immediately after stopping run 2892
    2894 05:40 05:58
    2895 05:59 06:23
    2896 06:23 06:46 detector reply ok
    2897 06:59 07:34 40uA on dummy target
    2898 07:47 08:00